Lauren plans to deposit $6000 into a bank account at the beginning of next month and $250/month into the same account at the end of that month and at the end of each subsequent month for the next 5 years. If her bank pays interest at a rate of 4%/year compounded monthly, how much will Lauren have in her account at the end of 5 years? (Assume she makes no withdrawals during the 5-year period. Round your answer to the nearest cent.)


1
Expert's answer
2021-03-29T11:03:25-0400

"FV=PV(1+r/n)^{nt-1}+A[(1+r/n)^{nt}-1]\div r/n"

"r=4\%, t=5yrs, n=12, A=250, PV=6000"

"FV=6000(1+0.04/12)^{(12)(5)-1}+250[(1+0.04/12)^{(12)(5)}-1]\div 0.04/12"

"6000\times1.2169+250\times 66.03=7301.4+16507.5=\$23,808.9"